# Optimized quantum sensor networks for ultralight dark matter detection

### v1

@@ Abstract
Introduction. Dark matter (DM), comprising about 27% of the universe’s mass-energy content, remains one of the most profound mysteries in modern physics [1, 2]. It is essential for explaining the gravitational binding of galaxies and the formation of large-scale cosmic structures [3, 4]. The concept originated from mass inconsistencies observed in galaxy clusters by Zwicky [5], and was later supported by galactic rotation curves measured by Rubin [6]. Although many experiments target potential DM candidates such as WIMPs, axions, and sterile neutrinos, none have been confirmed [7, 8, 9, 10, 11, 12, 13]. The detection of weakly interacting or light DM candidates, such as axions or hidden photons, remains challenging due to their extremely weak interactions with standard matter [14]. Identifying the fundamental nature of DM remains a central objective in both astrophysics and particle physics.
@@
Recent advances in quantum technologies have opened new avenues for DM detection. Superconducting qubits, in particular, offer exceptional sensitivity to weak external perturbations due to their tunability and long coherence times [15, 16]. It has been proposed that individual qubits can detect ultralight DM through shifts in electromagnetic fields or spin orientations [17], while entangling multiple qubits can enhance the sensitivity [18, 19]. Techniques from quantum metrology, which exploit quantum correlations to surpass classical measurement limits, have been applied to analyze the detection precision [20].
@@
Quantum sensor networks, formed by interconnected sensors, provide advanced capabilities for quantum sensing [21, 22]. By distributing quantum correlations across multiple nodes, such networks can improve both sensitivity and robustness, as demonstrated in star-shaped architectures [23, 24, 25]. However, network topology is crucial, as random or poorly structured configurations may significantly limit sensing performance. Network optimization is crucial for maximizing quantum sensing performance.
@@
In this work, we propose a network-based quantum sensor to enhance DM detection. Each node in the network is a superconducting qubit, connected via controlled-Z (CZ) gates. We explore symmetric network topologies, including line, ring, star, and fully connected graphs, to evaluate their impact on sensing performance. We analyze two representative systems comprising four and nine qubits, respectively.
@@
To optimize sensing performance, we employ a variational quantum metrology (VQM) approach [25, 26, 27], minimizing both the quantum Cramér-Rao bound (QB) and classical Cramér-Rao bound (CB) by variationally updating the initial quantum states and measurement settings. For estimation, we implement a Bayesian inference protocol to extract the DM–induced phase from measurement outcomes.
@@
Our results demonstrate that optimized network-based sensors outperform standard GHZ-based protocols in sensitivity. Optimal configurations achieve high performance with shallow circuits, making them suitable for noisy intermediate-scale quantum (NISQ) devices. The precision improves with increasing qubit number, consistent with theoretical expectations for quantum-enhanced sensing. It also remains robust against dephasing noise.
Beyond DM detection, the optimization methods and network architectures developed here are broadly applicable to a range of quantum sensing tasks, including gravitational wave detection, precision spectroscopy, and magnetic field sensing [28]. By integrating quantum computing, metrology, and fundamental physics, this approach provides a practical framework for addressing some of the most challenging problems in science.
@@
DM detection via superconducting qubits. Recent studies have used superconducting transmon qubits as quantum sensors for detecting ultralight DM [16, 17, 18, 19]. In these schemes, a coherently oscillating DM field induces an effective electric field, $\bar{E}^{(\text{eff})}$, which drives Rabi-like transitions between the ground and excited states of a transmon qubit. The interaction is described by the effective Hamiltonian
@@
where $\omega$ is the qubit energy splitting, $m_{\text{DM}}$ is the DM mass, $\alpha$ is the DM phase offset, and $|g\rangle$, $|e\rangle$ denote the ground and excited states, respectively. The coupling strength $\eta$ is given by
$\eta\equiv\frac{\sqrt{C\omega}}{2\sqrt{2}}\bar{E}^{(\text{eff})}d\cos\Theta,$ (2)
where $C$ and $d$ are the transmon’s capacitance and plate separation, and $\Theta$ is the angle between $\vec{E}^{(\text{eff})}$ and the normal vector of the capacitor plates. In simulations, $\cos^{2}\Theta\approx 1/3$ to account for angular fluctuations within the DM coherence time [18, 19].
The evolution of the qubit state is governed by the time-dependent Schrödinger equation
$i\frac{d}{dt}|\Psi(t)\rangle=\mathcal{H}|\Psi(t)\rangle,$ (3)
where $|\Psi(t)\rangle=\psi_{g}(t)|g\rangle+\psi_{e}(t)|e\rangle$ denotes the qubit wavefunction expressed in the ground and excited state basis. Within the coherence time $\tau\simeq\min(\tau_{\text{DM}},\tau_{q})$, where $\tau_{\text{DM}}$ and $\tau_{q}$ are the coherence times of the DM field and the qubit, respectively, the evolution is described by a unitary operator $U_{\text{DM}}(\tau)$
@@
In the resonant regime, $\omega=m_{\text{DM}}$, the interaction simplifies under the rotating wave approximation, and the evolution operator takes the form
@@
with eigenstates $|\psi_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|g\rangle\pm e^{i\alpha}|e\rangle\right).$
In the individual-qubit readout scheme, each transmon functions as an independent sensor, and the excitation probability induced by the DM field accumulates linearly with the number of qubits $p_{g\rightarrow e}(\tau)\simeq N\delta^{2}$, for $N\delta^{2}\ll 1,$ where $N$ is the number of qubits.
To surpass this shot-noise limit, quantum-enhanced sensing protocols based on entangled states such as GHZ states have been proposed [18, 19]. At the initial time $t_{1}$, a GHZ state is prepared
@@
During interacting with the DM field over a time interval $\tau=t_{2}-t_{1}$, each qubit acquires a phase shift $\delta$, resulting in the global state
@@
where |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|g\rangle\pm|e\rangle)$.
After the interaction, a reversed GHZ operation is applied to coherently map the accumulated phase information onto a single qubit
@@
Here, we assume $t_{f}=t_{2}$, neglecting further evolution beyond the sensing interval. The entangled state enables phase accumulation at a rate proportional to $N^{2}\delta^{2}$, yielding sensitivity that scales quadratically with $N$, in contrast to the linear scaling of individual-qubit strategies. This quadratic enhancement highlights a central advantage of entangled probes in quantum metrology, enabling parameter estimation beyond the standard quantum limit.
However, the optimal state for enhancing metrological precision is a superposition of the minimum and maximum eigenstates of the unitary evolution operator $U_{\rm DM}$, taking the form of a GHZ-like state [29]
$|\Psi(t_{1})\rangle=\frac{1}{\sqrt{2}}\Big{(}|\psi_{+}\rangle^{\otimes N}+|\psi_{-}\rangle^{\otimes N}\Big{)}.$ (9)
In contrast, the conventional GHZ state in Eq. (6) does not yield improved precision, as shown by the QB in Fig. 2(a). To overcome this limitation, we employ a network-based quantum sensor architecture to rigorously evaluate the achievable sensitivity. We use a VQM approach to optimize performance beyond what is attainable with GHZ states alone.
Figure 1: (a) A sensor-network architecture for detecting DM, which composes a 2D array of superconducting qubits connected via controlled-Z (CZ) gates to form a tunable interaction network. (b) Network configurations for $N=4$ qubits. (c) Network configurations for $N=9$ qubits. (d) Schematic of a VQM framework. The ansatze $U(\bm{\theta})$ (quantum state preparation) and $M(\bm{\mu})$ (measurement) are drawn from the structures in (b) and (c). The unitary $V(\delta,\alpha)$ encodes the external DM-induced evolution. The cost function $\mathcal{C}(\bm{\theta},\bm{\mu})$ guides the optimization based on QB and CB. The reconstruction is given through Bayesian estimation. (e) Example variational circuit structure for S4 configuration, showing the placement of single-qubit rotations and CZ gates.
Quantum-enhanced metrology viewpoint. The precision in estimating $\delta$ is fundamentally bounded by Cramér-Rao bounds
@@
where $\nu$ is the number of independent measurements, $F(\delta)$ is the classical Fisher information (CFI), and $Q(\delta)$ is the quantum Fisher information (QFI). These bounds quantify the ultimate limits on parameter estimation precision, with the QFI setting the best achievable sensitivity for any measurement strategy.
CFI depends on the measurement outcomes through the conditional probability distribution $p(m|\delta)$, and is given by
$F(\delta)=\sum_{m}\frac{1}{p(m|\delta)}\left(\frac{\partial p(m|\delta)}{\partial\delta}\right)^{2}.$ (11)
This expression quantifies the information about $\delta$ obtainable from a specific measurement strategy.
In contrast, QFI characterizes the maximum information attainable over all possible quantum measurements. It depends solely on the quantum state $\rho_{\delta}$ and is defined as
@@
where $\rho_{\delta}=|\Psi(t_{2})\rangle\langle\Psi(t_{2})|$ is the state after sensing and $L_{\delta}$ is the symmetric logarithmic derivative (SLD), determined by $2\partial\rho_{\delta}/\partial\delta=L_{\delta}\rho_{\delta}+\rho_{\delta}L_{\delta}$. A larger QFI indicates greater sensitivity, enabling more precise parameter estimation.
For separable (non-entangled) states, the QFI scales linearly with the number of qubits, $Q(\delta)\propto N,$ corresponding to the shot-noise limit, which represents the fundamental precision bound for classical strategies.
@@
In contrast, when using an entangled GHZ state, the QFI scales quadratically with the number of qubits, $Q(\delta)\propto 4N^{2},$ demonstrating Heisenberg-limited sensitivity. However, as noted above, standard GHZ states do not always yield optimal precision. This motivates further exploration of adaptive protocols and optimal measurement strategies, as discussed below.
Figure 2: (a) Comparison of QBs achieved by different ansatz configurations for a four-qubit system with $L_{1}=1$. The optimized ansatze, particularly the fully connected F4 architecture, outperform standard choices like the GHZ and excited states. (b) QB performance of F4 configuration as a function of the number of layers $L_{1}$. The best precision is achieved with a shallow circuit ($L_{1}=1$), highlighting the efficiency and practicality of the method for NISQ devices. (c) Robustness of the optimal configuration (F4, $L_{1}=1$) across a wide range of $\delta$ and $\alpha$. The QB remains nearly constant with $\delta$ and exhibits periodic variation with $\alpha$, demonstrating adaptability without the need for re-optimization.
@@
Quantum networks for DM detection. We consider a network-based quantum sensor modeled as a two-dimensional array of superconducting qubits, as illustrated in Fig. 1(a). Each qubit acts as a node, with edges representing physical connections between nodes. Due to hardware constraints, each qubit is limited to a maximum of four connections, consistent with the restricted connectivity in current superconducting architectures. We focus on symmetric network topologies known to enhance sensing precision. Specifically, we examine systems of four and nine qubits, shown in Fig. 1(b) and Fig. 1(c), respectively. For the 4-qubit case, we analyze linear (L4), star (S4), ring (R4), and fully connected (F4) configurations. For the 9-qubit system, we study analogous structures with linear (L9), star (S9), ring-star hybrid (RS9), and fully connected network (F9). These network architectures provide a structured platform for exploring the role of geometry in quantum sensing performance.
Variational quantum metrology. VQM is a hybrid quantum-classical framework for optimizing quantum sensing protocols using parameterized quantum circuits [25, 26, 27]. In this scheme, an ansatz $U(\bm{\theta})$ prepares the quantum state. Following interaction with the target process, such as the DM-induced unitary $U_{\rm DM}$, a second ansatz $M(\bm{\mu})$ is applied to optimize the measurement basis. These ansatze are governed by trainable parameter sets $\bm{\theta}$ and $\bm{\mu}$, respectively.
The optimization loop employs an optimizer to update $\bm{\theta}$ and $\bm{\mu}$ by minimizing (or maximizing) a cost function, which may be the QB, CB, or another task-specific metric. This hybrid framework enables adaptive learning of optimal sensing strategies and has been experimentally demonstrated on various quantum platforms [30, 31]. It is well suited for near-term devices that lack full error correction.
In this work, we implement the VQM scheme illustrated in Fig. 1(d), where the ansatze $U(\bm{\theta})$ and $M(\bm{\mu})$ are structured according to the network topologies shown in Figs. 1(b) and 1(c). These ansatze consist of rotation blocks $U_{\rm rot}=R_{z}(\theta_{z})R_{y}(\theta_{y})R_{x}(\theta_{x})$, inserted between each controlled-Z (CZ) gate, which defines the qubits connectivity [25]. An explicit example of the ansatz $U(\bm{\theta})$ for S4 configuration is provided in Fig. 1(e). We denote $L_{1}$ and $L_{2}$ as the number of layers in $U(\bm{\theta})$ and $M(\bm{\mu})$, respectively.
@@
Optimization and estimation. We begin by optimizing the parameters $\bm{\theta}$ by minimizing a cost function defined by the QB $\mathcal{C}(\bm{\theta})=Q^{-1}(\delta),$ yielding the optimal configuration $\bm{\theta}^{*}=\arg\min_{\{\bm{\theta}\}}\mathcal{C}(\bm{\theta})$. Once $\bm{\theta}^{*}$ is fixed, we optimize the measurement ansatz $M(\bm{\mu})$ using a similar procedure, with the cost function based on the CB, $\mathcal{C}(\bm{\mu})=F^{-1}(\delta).$ To update parameters, we use Adam optimizer.
After optimizing the network state and measurement configuration, we apply Bayesian estimation to infer the unknown parameter $\delta$ from the measurement outcomes. Let $m\in\{00\cdots 0,\ldots,11\cdots 1\}$ denote an outcome in the computational basis, and $p(\delta)$ the prior probability distribution representing initial knowledge of $\delta$. Here, we use the probability density function (pdf). The likelihood function $p(m|\delta)$ is given by the probability of observing outcome $m$ for a given $\delta$, computed from the final quantum state.
@@
[…]
@@
Quantum bound. We evaluate the performance of various ansatz configurations for a four-qubit system ($N=4$) and present the results in Fig. 2(a) for $L_{1}=1$. The GHZ and excited state (E) do not achieve QBs comparable to those of the optimized configurations. As noted earlier, the GHZ state is suboptimal under evolution by $U_{\rm DM}$ because it is not aligned with the eigenstates associated with the eigenvalues of the unitary. Consequently, its performance is comparable to that of the excited state.
@@
In Fig. 2(b), we explore the QB of F4 for different layers $L_{1}$. Remarkably, the best performance is achieved with just a single layer ($L_{1}=1$). This result is promising, as it implies that high precision can be obtained using shallow circuits, making the method both efficient and suitable for NISQ devices.
With the optimal configuration fixed at F4, $L_{1}=1$ and using the optimized parameters $\bm{\theta}$ found earlier, we next examine the QB as a function of the parameters $\delta$ and $\alpha$, as shown in Fig. 2(c). We observe that for a fixed $\alpha$, the QB remains constant over a wide range of $\delta$, from $10^{-3}$ to $10^{0}$. This robustness implies that re-optimization is not required for each new value of $\delta$, offering significant practical advantages for sensing the DM field. When varying $\alpha$, the QB shows periodic behavior, peaking at $\alpha=\pi/2$ and returning to its original value at $\alpha=0$ and $\pi$. Since the DM phase offset $\alpha$ is typically under experimental control, this does not negatively affect the practical performance.
@@
Figure 3: (a) CB as a function of the number of layers $L_{2}$ in the measurement circuit $M(\bm{\mu})$. Increasing $L_{2}$ improves the CB, bringing it closer to the QB, until an optimal point is reached. Beyond this, further increasing $L_{2}$ causes performance degradation. (b) CB versus different values of $\delta$ and $\alpha$. The optimized measurement configuration maintains a nearly constant and minimal CB across a wide range of $\delta$, except for relatively large $\delta$ where CB slightly increases. A periodic dependence on $\alpha$ is observed, consistent with the behavior seen in the QBs.
[…]
Figure 3(a) shows that with a shallow measurement circuit ($L_{2}=1$), the CB is larger than the QB, indicating poor measurement performance. As $L_{2}$ increases, the CB approaches the QB, reflecting improved accuracy. However, beyond an optimal depth, further increasing $L_{2}$ leads to a rise in the CB, reducing precision. This non-monotonic behavior suggests that deeper circuits may suffer from limited expressibility, overparameterization, or barren plateaus, where the optimization landscape becomes flat. The gap between CB and QB arises from the inability of the measurement ansatz to fully approximate the optimal POVM associated with the QFI.
In Fig. 3(b), we examine how the CB varies with different values of $\delta$ and $\alpha$. The results show that our optimized measurement configuration yields a constant, minimal CB over a wide range of $\delta$, similar to the behavior observed for the QB in Fig. 2(c). However, when $\delta$ becomes relatively large (on the order of $10^{0}$), the CB begins to increase. This rise does not significantly impact our conclusions, as the physically relevant regime for the DM field corresponds to small $\delta$ values.
Regarding $\alpha$, the phase offset of the DM interaction, we observe periodic variations in the CB, consistent with the behavior of the QB. This periodicity indicates that estimation performance depends on $\alpha$, but the optimal configuration remains robust across its range.
Bayesian estimation. Figure 4 presents the results of the Bayesian estimation procedure. In Fig. 4(a), we plot the posterior probability distribution as a function of $\delta$ for different numbers of measurements $\nu$. As $\nu$ increases, the peak of the posterior distribution shifts closer to the true value of $\delta$, demonstrating improved estimation accuracy.
@@
Figure 4: (a) Posterior probability distribution versus the phase $\delta$ for different number of measurements $\nu$. As $\nu$ increases, the posterior peak moves closer to the true value. (b) Bias $\bar{\delta}-\delta_{\rm true}$ as a function of $\nu$, showing convergence to zero with increasing $\nu$. (c) Variance Var($\delta$) versus $\nu$, compared with the ${\rm CB}/\nu$ and ${\rm QB}/\nu$. The results demonstrate good agreement, confirming the optimal scaling behavior.
@@
For $N=9$. We evaluate the QB and present the results in Fig. 5. Using the same procedure as in Fig. 2(a) for the four-qubit system, we examine multiple configurations shown in Fig. 1(c). Among these, the fully connected architecture F9 delivers the highest QB performance.
@@
When compared to the four-qubit case, the QB for $N=9$ is significantly improved. This improvement is expected because increasing the number of qubits generally enhances the achievable precision in quantum sensing, as predicted by quantum metrology theory.

### v2

@@ Abstract
Introduction. Dark matter (DM), comprising about 27% of the universe’s mass-energy content, remains a central mystery in modern physics [1, 2]. It is essential for explaining galactic binding and the formation of large-scale cosmic structures [3, 4]. The idea was first suggested by Zwicky’s discovery of missing mass in galaxy clusters [5] and later supported by Rubin’s measurements of galactic rotation curves [6]. Despite extensive searches for candidates such as WIMPs, axions, and sterile neutrinos, none have been confirmed [7, 8, 9, 10, 11, 12, 13]. Detecting weakly interacting or light candidates like axions or hidden photons remains difficult due to their extremely weak couplings with standard matter [14]. Revealing the fundamental nature of DM is still a central goal of astrophysics and particle physics.
@@
Recent advances in quantum technologies offer new opportunities for DM detection. Superconducting qubits provide exceptional sensitivity to weak perturbations owing to their tunability and long coherence times [15, 16]. Single qubits have been proposed to detect ultralight DM via shifts in electromagnetic fields or spin orientations [17, 18], while entangled qubits can further enhance sensitivity [19, 20]. Quantum metrology, exploiting correlations beyond classical limits, has been employed to analyze detection precision [21].
@@
Quantum sensor networks, built from interconnected sensors, enable enhanced sensitivity and robustness by distributing correlations across nodes [22, 23, 24, 25, 26]. However, their performance strongly depends on topology, e.g., poorly structured networks may degrade sensing. Network optimization is thus crucial for maximizing quantum sensing performance.
@@
In this Letter, we propose a quantum sensor network for enhanced DM detection, where each node is a superconducting qubit linked by controlled-Z (CZ) gates. We study symmetric topologies, i.e., line, ring, star, and fully connected. We analyze four- and nine-qubit systems to assess their sensing performance.
@@
To optimize sensing performance, we use variational quantum metrology (VQM) [26, 27, 28], which minimizes the quantum and classical Cramér-Rao bounds (QB, CB) by updating states and measurements. The DM-induced phase is then estimated via Bayesian inference.
@@
Our results show that optimized quantum sensor networks surpass GHZ-based protocols, achieving high precision with shallow circuits suitable for noisy intermediate-scale quantum (NISQ) devices. The precision scales favorably with qubit number and remains robust against dephasing noise.
Beyond DM detection, our optimization methods and network architectures extend to quantum sensing applications such as gravitational waves, spectroscopy, and magnetometry [29], providing a practical framework for tackling major scientific challenges.
@@
DM detection via superconducting qubits. Recent studies have used transmon qubits as quantum sensors for detecting ultralight DM [16, 17, 19, 20, 18]. In these schemes, a coherently oscillating DM field induces an effective electric field, $\bar{E}^{(\text{eff})}$, which drives Rabi-like transitions between the ground and excited states of the qubit. The interaction is described by the effective Hamiltonian
@@
where $\omega$ is the qubit energy splitting, $m_{\text{DM}}$ is the DM mass, $\alpha$ is the DM random phase offset, and $|g\rangle$, $|e\rangle$ denote the ground and excited states, respectively. The coupling strength is given by $\eta\equiv\frac{\sqrt{C\omega}}{2\sqrt{2}}\bar{E}^{(\text{eff})}d\cos\Theta,$ where $C$ is the transmon’s capacitance, $d$ is the distance between two plates, obeying $C\propto 1/d$, and $\Theta$ is the angle between $\vec{E}^{(\text{eff})}$ and the normal vector of the capacitor plates. Increasing $d$ decreases $C$ and increases the coupling $\eta$. For transmon qubits, $d\sim 10\text{–}100\,\mu\text{m}$. Fabrication limits, such as lithographic resolution, alignment, and chip size, further restrict $d$: too small $d$ is prone to defects, while very large $d$ increases chip area and parasitic effects. In simulations, $\cos^{2}\Theta\approx 1/3$ to account for angular fluctuations within the DM coherence time [19, 20].
The evolution is governed by the Schrödinger equation, $i\frac{d}{dt}|\Psi(t)\rangle=\mathcal{H}|\Psi(t)\rangle,$ where $|\Psi(t)\rangle=\psi_{g}(t)|g\rangle+\psi_{e}(t)|e\rangle$. Within the coherence time $\tau\simeq\min(\tau_{\text{DM}},\tau_{q})$, where $\tau_{\text{DM}}$ and $\tau_{q}$ are the coherence times of the DM field and the qubit, respectively, the evolution is described by a unitary operator $U_{\text{DM}}(\tau)$
@@
In the resonant regime, $\omega=m_{\text{DM}}$, the interaction simplifies under the rotating wave approximation, yielding
@@
where $|\psi_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|g\rangle\pm e^{i\alpha}|e\rangle\right).$ When $\omega\neq m_{\text{DM}}$, the sensitivity decreases over a mismatch bandwidth $\Delta f\sim\tau^{-1}$. For a fixed $\omega$, high sensitivity is maintained for $m_{\text{DM}}\in\omega\pm\tau^{-1}$, and $U_{\text{DM}}$ remains applicable [20].
In the individual-qubit scheme, each transmon acts as an independent sensor, giving a total excitation probability $p_{g\rightarrow e}(\tau)\simeq N\delta^{2}$, for $N\delta^{2}\ll 1,$ where $N$ the number of qubits.
To surpass this shot-noise limit, entangled protocols using GHZ states have been proposed [19, 20]. At $t_{1}$, a GHZ state is prepared
@@
During the interaction time $\tau=t_{2}-t_{1}$, each qubit acquires a phase $\delta$, giving
@@
where $|\pm\rangle=\frac{1}{\sqrt{2}}(|g\rangle\pm|e\rangle)$.
After interaction, a reversed GHZ operation maps the accumulated phase onto one qubit
@@
Here, we assumed $t_{f}=t_{2}$, neglecting further evolution beyond the sensing interval. The entangled state accumulates phase as $N^{2}\delta^{2}$, giving quadratic sensitivity in $N$, in contrast to the linear scaling of individual-qubit strategies. This quadratic enhancement highlights a central advantage of entangled probes in quantum metrology, enabling parameter estimation beyond the shot-noise limit.
However, the optimal state for enhancing metrological precision must be a superposition of the minimum and maximum eigenstates of $U_{\rm DM}$ [30], i.e, $|\Psi(t_{1})\rangle=\frac{1}{\sqrt{2}}(|\psi_{+}\rangle^{\otimes N}+|\psi_{-}\rangle^{\otimes N}).$ In contrast, the conventional GHZ state in Eq. (4) does not yield improved precision [see Fig. 2(a)]. To overcome this limitation, we employ a network-based quantum sensor and optimize performance using VQM.
Figure 1: (a) Sensor-network for DM detection: 2D superconducting qubit array with CZ gates forming a tunable network. (b)–(c) Configurations for $N=4$ and $N=9$. (d) VQM framework where $U(\bm{\theta})$ and measurement $M(\bm{\mu})$ come from (b, c); external evolution $V(\delta,\alpha)$ comes from the DM field; the cost $\mathcal{C}(\bm{\theta},\bm{\mu})$ optimized via QB and CB; reconstruction by Bayesian estimation. (e) Example variational circuit for S4.
Quantum-enhanced metrology viewpoint. The precision in estimating $\delta$ is bounded by Cramér-Rao bounds
@@
where $\nu$ is the number of measurements, $F(\delta)$ is the classical Fisher information (CFI), and $Q(\delta)$ is the quantum Fisher information (QFI).
The CFI, determined by the conditional probability $p(m|\delta)$, is
$F(\delta)=\sum_{m}\frac{1}{p(m|\delta)}\left(\frac{\partial p(m|\delta)}{\partial\delta}\right)^{2},$ (8)
quantifying the information about $\delta$ from a given measurement strategy.
The QFI gives the maximum information attainable from all quantum measurements and depends only on the state $\rho_{\delta}$
@@
where $\rho_{\delta}=|\Psi(t_{2})\rangle\langle\Psi(t_{2})|$ and $L_{\delta}$ is the symmetric logarithmic derivative defined by $2\partial\rho_{\delta}/\partial\delta=L_{\delta}\rho_{\delta}+\rho_{\delta}L_{\delta}$. Larger QFI implies higher sensitivity and more precise estimation.
@@
For non-entangled states, $Q(\delta)\propto N,$ corresponding to the shot-noise limit, which represents the fundamental precision bound for classical strategies. For an entangled GHZ state, $Q(\delta)\propto 4N^{2},$ demonstrating Heisenberg limit. However, as noted above, standard GHZ states do not always yield optimal precision. This motivates further exploration of adaptive protocols and optimal measurement strategies, as discussed below.
@@
Quantum networks for DM detection. We model a network-based quantum sensor as a 2D array of superconducting qubits, as illustrated in Fig. 1(a). Each qubit is a node, and edges represent physical connections between nodes. Due to hardware constraints, each qubit is limited to a maximum of four connections, consistent with the restricted connectivity in current superconducting architectures. We examine systems of four and nine qubits, shown in Fig. 1(b) and Fig. 1(c), respectively. For the 4-qubit case, we analyze linear (L4), star (S4), ring (R4), and fully connected (F4) configurations. For the 9-qubit system, we study simular structures with linear (L9), star (S9), ring-star hybrid (RS9), and fully connected network (F9).
Variational quantum metrology. VQM is a hybrid quantum-classical framework for optimizing quantum sensing protocols using parameterized quantum circuits [26, 27, 28]. In this scheme, the state is prepared by an ansatz $U(\bm{\theta})$, evolves under an external field such as $U_{\rm DM}$, and is then measured through an ansatz $M(\bm{\mu})$. The parameters $\bm{\theta}$ and $\bm{\mu}$ are variationally updated to minimize cost functions such as the Cramér-Rao bounds [26]. This hybrid framework enables adaptive sensing strategies, has been demonstrated on several platforms [31, 32], and is well suited to near-term devices.
VQM shares a structure with other variational algorithms like variational quantum eigensolver (VQE) [33], where a parameterized quantum circuit is optimized via classical feedback. While VQE minimizes energy to find ground states, VQM minimizes the Cramér-Rao bounds to improve the parameter estimation, highlighting their similar workflows.
Figure 2: (a) QBs for four-qubit system at $L_{1}=1$; optimized ansatze, especially F4, outperform GHZ and excited states. (b) QBs of F4 versus depth $L_{1}$; best precision at shallow $L_{1}=1$, showing efficiency. (c) Robustness of F4 at $L_{1}=1$; QB nearly constant with $\delta$ and periodic in $\alpha$ without re-optimization.
We implement the VQM illustrated in Fig. 1(d), where $U(\bm{\theta})$ and $M(\bm{\mu})$ are structured according to the network in Figs. 1(b, c). These ansatze consist of rotation blocks $U_{\rm rot}=R_{z}(\theta_{z})R_{y}(\theta_{y})R_{x}(\theta_{x})$, inserted between each CZ gate [26]. An example of $U(\bm{\theta})$ for S4 is provided in Fig. 1(e). We denote $L_{1}$ and $L_{2}$ as the number of layers in $U(\bm{\theta})$ and $M(\bm{\mu})$, respectively. These ansatze use a linear number of parameters, keeping the optimization cost low for small systems. In larger Hilbert spaces, shallow circuits may lack expressivity [27, 28], whereas deeper circuits improve it but introduce risks such as barren plateaus [34], initialization sensitivity, and increased noise susceptibility. Therefore, selecting the optimal depth involves balancing expressivity, trainability, and robustness.
@@
Optimization and estimation. We first optimize $\bm{\theta}$ by minimizing a cost function defined by $\mathcal{C}(\bm{\theta})=Q^{-1}(\delta),$ yielding the optimal configuration $\bm{\theta}^{*}=\arg\min_{\{\bm{\theta}\}}\mathcal{C}(\bm{\theta})$. Once $\bm{\theta}^{*}$ is fixed, we optimize $\bm{\mu}$ using a similar procedure, with the cost function is $\mathcal{C}(\bm{\mu})=F^{-1}(\delta).$ To update parameters, we use Adam optimizer.
After optimizing the network state and measurement configuration, we apply Bayesian estimation to recover $\delta$ from the measurement outcomes. Let $m\in\{00\cdots 0,\ldots,11\cdots 1\}$ denote an outcome in the computational basis, and $p(\delta)$ the prior probability distribution representing initial knowledge of $\delta$. The likelihood function $p(m|\delta)$ is given by the probability of observing outcome $m$ for a given $\delta$, computed from the final quantum state.
@@
where $p(m)=\int p(m|\delta)p(\delta)\,d\delta$ is the marginal likelihood. We choose $p(\delta)$ as the probability density function (pdf). This choice has minimal impact on the final results because each measurement updates the posterior, which then serves as the new prior. For a large number of measurements, this sequential updating ensures the final estimate is largely insensitive to the initial prior.
@@
Quantum bound. We evaluate the performance of various ansatze for a four-qubit system ($N=4$) and present the results in Fig. 2(a) for $L_{1}=1$. The GHZ and excited state (E) do not achieve QBs comparable to those of the optimized configurations. As noted earlier, the GHZ state is suboptimal under evolution by $U_{\rm DM}$ because it is not aligned with the eigenstates associated with the eigenvalues of the unitary. Consequently, its performance is comparable to that of the excited state.
@@
In Fig. 2(b), we explore the QB of F4 for different layers $L_{1}$. Remarkably, the best performance is achieved with $L_{1}=1$. The result indicates that high precision is achievable with shallow circuits, highlighting the method’s efficiency and practicality for NISQ devices. Nonetheless, practical limitations, including gate errors ($\sim$ 0.1–1%), finite coherence times, and restricted connectivity in superconducting devices, can affect performance. A detailed analysis will be addressed in future work.
With the fixed F4, $L_{1}=1$ configuration, we next examine the QB as a function of $\delta$ and $\alpha$, as shown in Fig. 2(c). For a fixed $\alpha$, the QB remains constant over a wide range of $\delta$, from $10^{-3}$ to $10^{0}$. This robustness implies that re-optimization is not required for each new value of $\delta$, offering significant practical advantages for sensing the DM field. When varying $\alpha$, the QB shows a periodic behavior, peaking at $\alpha=\pi/2$ and returning to its original value at $\alpha=0$ and $\pi$. While the optimization was performed at $\alpha=0$, we verified that the protocol performs robustly for all $\alpha$ values. The QB shows only an increase around $\alpha=\pi/2$, which does not significantly affect the overall measurement precision.
@@
Classical bound. To achieve the best CB, it is essential to use the optimal measurement operators, represented as positive operator-valued measures (POVMs). In our VQM framework, these measurement operators are optimized through the ansatz $M(\bm{\mu})$. Specifically, the circuit for $M(\bm{\mu})$ is chosen to be the inverse of the input circuit, i.e., ${\rm F}4^{\dagger}$, with a variable number of layers $L_{2}$.
Figure 3(a) shows that with a shallow measurement circuit ($L_{2}=1$), the CB is larger than the QB, indicating poor measurement performance. As $L_{2}$ increases, the CB approaches the QB, reflecting improved accuracy. However, beyond an optimal depth, further increasing $L_{2}$ will raising the CB and reducing precision. This non-monotonic behavior suggests that deeper circuits may suffer from limited expressibility, overparameterization, or barren plateaus, where the optimization landscape becomes flat. The gap between CB and QB arises from the inability of the measurement ansatz to fully approximate the optimal POVM associated with the QFI.
In Fig. 3(b), we examine how the CB varies with $\delta$ and $\alpha$. Our optimized measurement configuration yields a constant, minimal CB over a wide range of $\delta$, similar to the behavior observed for the QB in Fig. 2(c). However, when $\delta$ becomes relatively large (on the order of $10^{0}$), the CB begins to increase. This rise does not significantly impact our conclusions, as the physically relevant regime for the DM field corresponds to small $\delta$ values.
Regarding $\alpha$, we observe periodic variations in the CB, consistent with the behavior of the QB. This periodicity indicates that estimation performance depends on $\alpha$, but the optimal configuration remains robust across its range.
Figure 3: (a) CB versus depth $L_{2}$; increasing $L_{2}$ decreases CB toward the QB, then increases. (b) CB versus $\delta$ and $\alpha$; optimized measurements keep CB nearly constant, with slight growth at large $\delta$ and periodic variation with $\alpha$, consistent with QBs.
Bayesian estimation. In Fig. 4(a), we plot the posterior probability distribution as a function of $\delta$ for different numbers of measurements $\nu$. As $\nu$ increases, the peak shifts closer to the true value of $\delta$, demonstrating improved estimation accuracy.
@@
For $N=9$. We evaluate the QB and present the results in Fig. 5. Using the same procedure as in Fig. 2(a) for the four-qubit system, we examine multiple configurations shown in Fig. 1(c). Among these, the fully connected F9 delivers the highest QB performance.
@@
When compared to $N=4$ , the QB for $N=9$ is significantly improved. This improvement is expected because increasing the number of qubits generally enhances the achievable precision in quantum sensing, as predicted by quantum metrology theory.